# Super approximation for $\text{SL}_{2}\times\text{SL}_{2}$ and $\text{ASL}_{2}$

### v3

@@ 6.1. The case $q_{4}^{\prime}>q_{4}^{\frac{1}{2}}$
We divide our proof into two cases.
7.1. The case $q_{s}>q^{\frac{\epsilon}{2}}$
Take
$\delta<\frac{\log\frac{1}{c_{0}}}{2}$ (7.2)
for $c_{0}=c_{0}(L)$ the implied constant from Theorem A, so that
$|\pi_{q_{s}}^{*}(\chi_{S}^{(l)})(x)-\frac{1}{|\Gamma_{q_{s}}|}|<\frac{1}{2}$
@@ 7.1. The case $q_{s}>q^{\frac{\epsilon}{2}}$
$|\pi_{q_{s}}(A)|>|\Gamma_{q_{s}}|q^{-\delta}>q_{s}^{6-\frac{2\delta}{\epsilon}-}>q_{s}^{6-2\rho}$ (7.3)
@@
$\delta<\min\{\frac{\log\frac{1}{c_{0}}}{2},{\epsilon\rho}\}.$ (7.4)
for some $\rho\ll\epsilon$ to be determined at (7.9).
Furthermore, if $q_{s}>q^{1-\frac{\epsilon}{2}}$, then (7.3) implies the assumption in (2.3) is void and so Proposition 2.2 automatically holds. So now we assume $q_{s}<q^{1-\epsilon/2}$.
Now we apply Proposition 5.1 with $q$ replaced by $q_{l}$, then we obtain some $q^{\prime}||q_{l},q^{\prime}>q_{l}^{c}>q^{\frac{\epsilon c}{2}}$ for some $c$ only depending on the generating set $S$, and $C^{\prime}=O_{S}(\log\frac{1}{\rho})$ such that
$|\pi_{q^{\prime}}\circ\mathbb{P}_{1}(A^{C^{\prime}})|>(q^{\prime})^{3-\rho}$
if
$\delta<c^{\prime}\epsilon\rho$ (7.5)
for some $c^{\prime}$ depending only on $S$.
Now we apply Proposition 6.1 with $B=A^{C^{\prime}}$, $q_{1}=q_{2}=q_{s}$, $q_{3}=q^{\prime}$ and $\theta=\rho$. We then obtain $q^{\prime\prime}|q^{\prime},q^{\prime\prime}>(q^{\prime})^{\frac{1}{4}\times 10^{-4}}$, and
$|\pi_{q_{s}q^{\prime\prime},q_{s}}(A^{C^{\prime}C^{\prime\prime}})|>(q_{s}^{2}q^{\prime\prime})^{3-O(\theta^{\frac{1}{4}})},$ (7.6)
for some
$C^{\prime\prime}=O(\log{1}/{\rho}).$ (7.7)
By (7.7), we have increased the modulus of the first component from $q_{s}$ to $q_{s}q^{\prime\prime}$, at the cost of a density loss from $3-\rho$ to $3-O(\rho^{{1}/{4}})$. If $q_{s}q^{\prime\prime}<q^{1-\frac{\epsilon}{2}}$, we apply Proposition 5.1 again with $q$ replaced by $\frac{q}{q_{s}q^{\prime\prime}}$ and $\theta=O(\rho^{1/4})$, after that we can apply Proposition 6.1 again to increase the modulus to a larger one. We apply Proposition 5.1 and Proposition 6.1 iteratively until we reach a modulus $q_{1}^{\star}||q,q_{1}^{\star}>q^{1-\frac{\epsilon}{2}}$. Next, we go through the same procedure to increase the modulus of the second component to $q_{2}^{\star}||q,q_{2}^{\star}>q^{1-\frac{\epsilon}{2}}$. The number of such iterations is bounded by by $\frac{K_{0}\log{(1/\epsilon)}}{c}$ for some absolute constant $K_{0}$.
In the end, we obtain
$\left|\pi_{q_{1}^{\star},q_{2}^{\star}}\left(A^{O_{S}(\log\frac{1}{\rho})}\right)\right|>(q_{1}^{\star}q_{2}^{\star})^{3-O(\rho^{(\frac{1}{4})^{\frac{K_{0}\log(1/\epsilon)}{c}}})}.$ (7.8)
Let $K_{1},K_{2}$ be the implied constants on the LHS and RHS of (7.8), respectively. If we take
$\rho=e^{-(\log(3K_{2}/\epsilon)4^{{\frac{K_{0}\log(1/\epsilon)}{c}}})},$ (7.9)
then from (7.8),
$\left|\pi_{q_{1},q_{2}}\left(A^{K_{1}(\log\frac{1}{\rho})}\right)\right|>q^{6-5\epsilon}.$ (7.10)
Recall also $\pi_{q}(A)<q^{6-6\epsilon}$. Then (7.10) and (5.13) imply
$|\pi_{q}(A\cdot A\cdot A)|>|A|^{1+\frac{\epsilon}{6K_{1}\log(1/\rho)}}.$ (7.11)
We take any $\delta_{0}$ sufficiently small satisfying (7.2), (7.4) and (7.5) for $\delta$. Then we take
$L=\left[\frac{13}{\delta_{0}}\right]$ (7.12)
in view of the condition (6.3). Finally, we set
$\delta=\min\{\delta_{0},\frac{\epsilon}{6K_{1}\log(1/\rho)}\}.$
Then (7.11) contradicts (7.1). Proposition 2.2 is thus proved in this case.
7.2. The case $q_{s}<q^{\frac{\epsilon}{2}}$
In this case the modulus $q_{s}$ can be ignored. We can directly apply Proposition 5.1 and start the iterating process. The analysis is identical to the previous case.
@@ 7.2. The case $q_{s}<q^{\frac{\epsilon}{2}}$
We write $q=q_{s}q_{l}$, where $q=\prod_{i\in I}p_{i}^{n_{i}}$, $q_{s}=\prod_{i\in I:n_{i}\leq L}p_{i}^{n_{i}}$, $q_{l}=\prod_{i\in I:n_{i}>L}p_{i}^{n_{i}}$ for some $L$ depending only on the generating set $S$ and $\epsilon$.
We take $\delta$ sufficiently small and assume
$|\pi_{q}(A\cdot A\cdot A)|<|\pi_{q}(A)|^{1+\delta}.$ (8.1)
If $q_{s}>q^{1-\epsilon/2}$, $q_{l}$ can be ignored and we can apply Theorem A to conclude Proposition 2.2.
If $q^{\epsilon/2}<q_{s}<q^{1-\epsilon/2}$, we first apply Theorem A to find some $C_{1}$ so that $A^{C_{1}}(\operatorname{mod}q_{s})$ is large in $\Lambda/\Lambda(q_{s})$. Then we use the method in Section 7 to find some $C_{2}$ so that $\pi_{q,q_{s}}(A^{C_{2}})$ is large in $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\ltimes(\mathbb{Z}/q_{s}\mathbb{Z})^{2}$, and so by Proposition 3.20, there exists $q_{1}||q,q_{2}||(q_{1},q_{s}),q_{1}q_{2}>(qq_{s})^{1-\epsilon/2}$ such that
$\pi_{q_{1},q_{2}}(A^{C_{2}})\supset(\Gamma(q_{1}^{\{\rho\}})/\Gamma(q_{1}))\ltimes(q_{2}^{\{\rho\}}\mathbb{Z}/q_{2}\mathbb{Z})$ (8.2)
for some $\rho=O(\delta)$. Choose $B\subset A^{C_{2}}$ a representative set of RHS of (8.2) and consider the section map
$\psi:(\Gamma(q_{1}^{\{\rho\}})/\Gamma(q_{1}))\ltimes(q_{2}^{\{\rho\}}\mathbb{Z}/q_{2}\mathbb{Z})\rightarrow B.$
If for many $p^{n}||\text{gcd}(q_{l},q_{1})$, $\pi_{p^{\{n\rho^{1/2}\}}}\circ\psi$ is not a homomorphism, then one can apply the argument in Section 6.1 find a large divisor $q_{3}||\text{gcd}(q_{l},q_{1})$ and an element $(g_{0},w_{0})$ from a bounded product of $B$ such that $\pi_{q_{1},q_{2}}((g_{0},w_{0}))=(1,0)$, and $v_{p}(w_{0})=O(\rho n^{1/2})$ for any $p^{n}||q_{3}$. Conjugating $(g_{0},w_{0})$ by $B$ using the formula $(g,x)\cdot(1,y)\cdot(g,x)^{-1}=(1,g(y)),$ and multiplying these conjugates by $B$, we obtain a set which is large in $\Lambda/\Lambda(q_{1})\ltimes(\mathbb{Z}/q_{2}q_{3})^{2}$ under the reduction $\pi_{q_{1},q_{2}q_{3}}$.
If for many $p^{n}||\text{gcd}(q_{l},q_{1})$, $\pi_{p^{\{n\rho^{1/2}\}}}\circ\psi$ is a homomorphism, then we can find a large divisor $q_{4}||\text{gcd}(q_{l},q_{1})$ and a large subset $Z\subset(\Gamma(q_{1}^{\{\rho\}})/\Gamma(q_{1}))\ltimes(q_{2}^{\{\rho\}}\mathbb{Z}/q_{2}\mathbb{Z})$ such that $\psi|_{Z}$ is a homomorphism. But then Proposition 3.24 implies $\psi(Z)(\operatorname{mod}q_{4}^{\{\Theta(\rho^{1/2})\}})\subset H(\operatorname{mod}q_{4}^{\{\Theta(\rho^{1/2})\}})$ for some proper algebraic subgroup $H\subset\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ defined by two linear equations. Then using Proposition 4.4 and arguing as in Section 5.1 for creating the set $A_{3}$, we can find a large exact divisor $q_{5}||q_{4}$ and $(g_{1},w_{1})$ from a bounded product of $B\cup(A\cdot A)$, such that $\pi_{q_{1},q_{2}}((g_{1},w_{1}))=(1,0)$, and $v_{p}(w_{1})=O(\rho^{1/2}n)$ for any $p^{n}||q_{5}$. With the help of $(g_{1},w_{1})$, we can boundedly generate a set which is large in $\text{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})\ltimes(\mathbb{Z}/q_{2}q_{5}\mathbb{Z})^{2}$.
In either of the above two cases, we increase the modulus $q_{2}$ by a factor of size at least $q^{\kappa}$ where $\kappa>0$ depending only on $S$ and $\epsilon$, Iterating the above procedure by controlled number of times, we can grow $q_{2}$ to a modulus $q_{6}||q_{1}$ and $q_{6}>q_{1}^{1-\epsilon/2}$, which will lead to a contradiction to (8.1) when $\delta$ is sufficiently small.
If $q_{s}<q^{\epsilon/2}$, then $q_{s}$ can be ignored and the expanding procedure is virtually identical to the previous case.

### v4

@@ 6.1. The case $q^{J_{1}}\geq(q_{3}^{\prime})^{\frac{1}{2}}$
We first observe if the first two conditions of (2.3) and (7.1) hold, then they also hold with $q$ replaced by any exact divisor $q^{\prime}$ of $q$ such that $q^{\prime}>q^{\frac{\varepsilon}{2}}$, and $\delta$ replaced by any
$\delta_{0}\geq\frac{48\delta}{\varepsilon c_{1}}.$ (7.2)
Indeed, (7.1) implies a similar product bound with modulus $q^{\prime}$. Divide $\pi_{q}(A)$ into congruence classes mod $q^{\prime}$, then at least one congruence class has cardinality $\geq\frac{|\pi_{q}(A)|}{|\pi_{q^{\prime}}(A)|}$. Recall by Proposition 4.1, $|\pi_{q^{\prime}}(A)|>(q^{\prime})^{\frac{c_{1}}{2}}$ if
$\delta\leq\frac{c_{1}}{2}.$ (7.3)
Therefore,
$|\pi_{q^{\prime}}(A)\cdot\pi_{q^{\prime}}(A)\cdot\pi_{q^{\prime}}(A)|\cdot\frac{|\pi_{q}(A)|}{|\pi_{q^{\prime}}(A)|}\leq|\pi_{q}(A)\cdot\pi_{q}(A)\cdot\pi_{q}(A)\cdot\pi_{q}(A)|\leq|\pi_{q}(A)|^{1+2\delta}$
$\Rightarrow$ $\frac{|\pi_{q^{\prime}}(A)\cdot\pi_{q^{\prime}}(A)\cdot\pi_{q^{\prime}}(A)|}{|\pi_{q^{\prime}}(A)|}\leq|\pi_{q}(A)|^{2\delta}<q^{12\delta}<(q^{\prime})^{\frac{24\delta}{\varepsilon}}<|\pi_{q^{\prime}}(A)|^{\frac{48\delta}{\varepsilon c_{1}}}\leq|\pi_{q^{\prime}}(A)|^{\delta_{0}}.$
We take $\delta_{0}$ small enough so it satisfies the “sufficiently small” requirement for $\delta$ at Proposition 5.1. A proper choice for $\delta_{0}$ is made at (7.8). Set
$L=\frac{3}{\delta_{0}},$ (7.4)
so that Proposition 5.1 is applicable with $q$ replaced by all exact divisors $q^{\prime}>q^{\frac{\varepsilon}{2}}$ and $\delta$ replaced by $\delta_{0}$. We denote the constants $\rho_{0}=\rho(\delta_{0}),C_{0}=C(\delta_{0}),c_{0}=c$ implied by Proposition 5.1.
We first apply Proposition 5.1 to obtain $q_{1}\|q,q_{1}\geq q^{c_{0}}$ such that
$\Gamma(q_{1}^{\{\rho_{0}\}})/\Gamma(q_{1})\subset\mathbb{P}_{1}(A)^{C_{0}}(\operatorname{mod}q_{1}).$
If $q_{1}>q^{1-\frac{\varepsilon}{2}}$, we are done. Otherwise, apply Proposition 5.1 again, with $q$ replaced by $\frac{q}{q_{1}}$, we obtain $q_{2}^{\prime}\|\frac{q}{q_{1}}$, $q_{2}^{\prime}\geq(\frac{q}{q_{1}})^{c_{0}}\geq q^{\frac{\varepsilon c_{0}}{2}}$, such that
$\Gamma((q_{2}^{\prime})^{\{\rho_{0}\}})/\Gamma(q_{2}^{\prime})\subset\mathbb{P}_{1}(A)^{C_{0}}(\operatorname{mod}q_{2}^{\prime}).$
We will take
$\delta\leq c_{0}\epsilon\rho_{0},$ (7.5)
so we can apply Proposition 6.1 to obtain $q_{2}^{\prime\prime}\|q_{2}$, $q_{2}^{\prime\prime}>(q_{2}^{\prime})^{\frac{1}{4}10^{-4}}$, such that
$\left|\pi_{q_{1}q_{2}^{\prime\prime}}\circ\mathbb{P}_{1}(A^{C_{0}K_{1}[\log\frac{1}{\rho_{0}}]^{2}})\right|>(q_{1}q_{2}^{\prime\prime})^{3-K_{2}\rho_{0}^{\frac{1}{4}}}$
for some two absolute constants $K_{1},K_{2}>0$ implied by the two big O notations in (6.4).
Write $q_{2}=q_{1}q_{2}^{\prime\prime}$. If $q_{2}>q^{1-\frac{\varepsilon}{2}}$, we are done. Otherwise, we apply Proposition 5.1 and Proposition 6.1 again to find $q_{2}\|q_{3}\|q$, with $\frac{q_{3}}{q_{2}}>q^{\frac{10^{-4}\varepsilon c_{0}}{8}}$, such that
$\left|\pi_{q_{3}}\circ\mathbb{P}_{1}(A^{C_{0}K_{1}^{2}[\log\frac{1}{\rho_{0}}]^{4}})\right|>q_{3}^{3-K_{2}^{\frac{5}{4}}\rho_{0}^{\frac{1}{16}}},$
and we keep going, until we find an integer $q_{T}\|q$, $q_{T}>q^{1-\frac{\varepsilon}{2}}$, with
$\left|\pi_{q_{T}}\circ\mathbb{P}_{1}(A^{C_{0}K_{1}^{T}[\log\frac{1}{\rho_{0}}]^{2T}})\right|=|\pi_{q_{T},1}(\cdots)|>q_{T}^{3-K_{2}^{2}\rho_{0}^{(\frac{1}{4})^{T}}}.$ (7.6)
The number of iterations $T$ is bounded by $\frac{8\cdot 10^{4}}{\varepsilon c_{0}}$.
Now we apply Proposition 5.1 and Proposition 6.1 to the $A$-power at 7.6 to expand on the second modulus. Iterate for another $T^{\prime}\leq\frac{8\cdot 10^{4}}{\varepsilon c_{0}}$ times, we obtain an integer $q_{T^{\prime}}^{\prime}\|q$, $q_{T^{\prime}}^{\prime}>q^{1-\frac{\varepsilon}{2}}$, such that
$\left|\pi_{q_{T},q_{T^{\prime}}^{\prime}}(A^{C_{0}K_{1}^{T+T^{\prime}}[\log\frac{1}{\rho_{0}}]^{2(T+T^{\prime})}})\right|>(q_{T}q_{T^{\prime}}^{\prime})^{3-K_{2}^{2}\rho_{0}^{(\frac{1}{4})^{T+T^{\prime}}}}.$ (7.7)
The exponent for $A$ at (7.7) is upper bounded by
$T_{0}:=C_{0}K_{1}^{\frac{16\cdot 10^{4}}{\varepsilon c_{0}}}[\log\frac{1}{\rho_{0}}]^{\frac{16\cdot 10^{4}}{\varepsilon c_{0}}},$
and the exponent for $q_{T}q_{T^{\prime}}^{\prime}$ is lower bounded by
$3-K_{2}^{2}\rho_{0}^{[(\frac{1}{4})^{\frac{16\cdot 10^{4}}{\varepsilon c_{0}}}]}.$
We take $\delta_{0}$ small enough so that $\delta_{0}$ satisfies the “sufficiently small” requirement for $\delta$ at Proposition 5.1 and the implied constant $\rho_{0}=\rho_{0}(\delta_{1})=O(\delta_{1})$ satisfies
$K_{2}^{2}\rho_{0}^{[(\frac{1}{4})^{\frac{16\cdot 10^{4}}{\varepsilon c_{0}}}]}<\frac{\varepsilon}{2}.$ (7.8)
This determines $T_{0}$ and $L$ subsequently.
(7.7) then implies
$|\pi_{q}(A)|^{T_{0}}>q^{6-5\varepsilon}>q^{\varepsilon}|\pi_{q}(A)|>|\pi_{q}(A)|^{1+\frac{\varepsilon}{6}}$
[…]
Thus recalling (7.2), (7.3) and (7.5), if we set
$\delta=\min\{\frac{\varepsilon c_{1}\delta_{0}}{48},\varepsilon c_{0}\rho_{0},\frac{c_{1}}{2},\frac{\varepsilon}{6T_{0}}\},$
then (7) contradicts (7.1). So (2.4) has to hold for the above choice of $\delta$. We have thus proved Proposition 2.2 for all sufficiently large $q$ with exponents of all prime divisors $\geq L$.
For a general $q$, let $q=q_{s}q_{l}$, where
$q_{s}=\prod_{i\in I,n_{i}\leq L}p_{i}^{n_{i}},\hskip 28.45274ptq_{l}=\prod_{i\in I:n_{i}>L}p_{i}^{n_{i}}.$
If $q_{s}\leq q^{\frac{\varepsilon}{2}}$, then the modulus $q_{s}$ can be ignored. We can work with $q_{l}$ and run the previous argument.
If $q_{s}>q^{\frac{\varepsilon}{2}}$, take
$\delta<\frac{c_{s}}{4}$ (7.10)
for $c_{s}=c_{s}(L)$ the implied spectral gap from Theorem A for $\Lambda$ and for all integer moduli with exponents of prime divisors $\leq L$. Then
$|\pi_{q_{s}}^{*}(\chi_{S}^{(l)})(x)-\frac{1}{|\Gamma_{q_{s}}|}|\leq\frac{1}{2|\Gamma_{q_{s}}|}$
@@
$|\pi_{q_{s}}(A)|>\frac{|\Gamma_{q_{s}}|}{2}q^{-\delta}>q_{s}^{6-\frac{2\delta}{\varepsilon}-}>q_{s}^{6-2\rho_{0}}$ (7.11)
@@
$\delta<\min\{\frac{c_{s}}{4},\frac{{\varepsilon\rho_{0}}}{2}\}.$ (7.12)
where $\rho_{0}$ is given after (7.4).
In case $q_{s}>q^{1-\frac{\varepsilon}{2}}$, (7.11) implies the third assumption in (2.3) is void and so Proposition 2.2 automatically holds.
In case $q_{s}\leq q^{1-\frac{\varepsilon}{2}}$, we can run the argument in the large exponent case to grow the modulus $(q_{s},q_{s})$ to $(q_{1}^{*},q_{2}^{*})$, where $q_{1}^{*},q_{2}^{*}>q^{1-\frac{\varepsilon}{2}}$ and the projection of product set of $A$ to $(q_{1}^{*},q_{2}^{*})$ is very large.
@@
Let $\mathbb{P}_{0}:\Lambda\rightarrow\Gamma$ be the projection to the semisimple part and $\mathbb{P}_{u}:\Lambda\rightarrow\mathbb{Z}^{2}$ the projection to the unipotent part.
We assume $q=\prod_{i\in I}p_{i}^{n_{i}}$ with each $n_{i}$ large. Assume (2.3) holds but (2.4) fails. Then by taking $\delta$ sufficiently small, we can apply Proposition 5.1 and Proposition 6.1 iteratively to obtain
$G:=\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\subset\mathbb{P}_{0}(A^{O((\log\frac{1}{\rho})^{2})}),$ (8.1)
where $\mathbb{P}_{0}$ is the projection to the semisimple part, $\rho$ is a small quantity, $q_{1}=q^{I_{1}}$ for some $I_{1}\subset I$ with $q_{1}>q^{1-\frac{\varepsilon}{2}}$, and for any $i\in I_{1}$, $v_{p_{i}}(q_{1}^{\prime})<\rho v_{p_{i}}(q_{1})$.
Let $B\subset A^{O((\log\frac{1}{\rho})^{2})}$ be a set of representatives of $G$ implied by (8.1). Let $\psi:G\rightarrow B$ such that $\mathbb{P}_{0}\circ\psi:G\rightarrow G$ is the identity map.
For each $i\in I_{1}$, consider $\psi_{i}=\pi_{p_{i}^{\{\theta n_{i}\}}}\circ\mathbb{P}_{u}\circ\psi$ for another small quantity $\theta$ which is larger than $\rho$.
According to Proposition 3.1, there are two scenarios:
Event 1: Define
$\mathcal{G}_{j}=\{(x,y)\in G\times G|\psi_{i}(xy)\neq\psi_{i}(x)\psi_{i}(y)\}$
We have,
$|\mathcal{G}_{j}|>10^{-4}|G|^{2}.$ (8.2)
Event 2: There is a subset $S_{i}\in G$, $|S_{i}|\geq\frac{99}{100}|G|$ such that $\psi_{i}\equiv h_{i}$ over $S_{i}$}, where $h_{i}$ is a homomorphism from $G$ to $\Gamma/\Gamma(p_{i}^{[n_{i}\theta]})$.
We split $I_{1}=I_{2}\sqcup I_{3}$, where $I_{2}$ consists of all $i$ such that Event 1 happens, and $I_{3}$ is the complement of $I_{2}$. We divide our analysis into two cases according to whether $q^{I_{2}}$ is large or not.
Case 1: $q^{I_{2}}\geq(q^{I_{1}})^{\frac{1}{2}}$. We can find $(1,w)\in B^{2}B^{-1}$ such that $w\in\mathbb{Z}^{2}$ satisfies $v_{p_{i}}(w)<\theta n_{i}$ for all $i\in I_{2}$. Conjugate $(1,w)$ by $B$ using the formula
$(g,x)\cdot(1,w)\cdot(g,x)^{-1}=(1,g(w)),$
we obtain a set $B_{1}$ such that $\mathbb{P}_{0}(B_{1})=1$ and $\mathbb{P}_{u}(B_{1})(\operatorname{mod}q^{I_{2}})$ is large in $(\mathbb{Z}/q^{I_{2}}\mathbb{Z})^{2}$. Therefore, $\pi_{q,q^{I_{2}}}(B\cdot B_{1})$ is a large subset of $\Gamma_{q}\ltimes(\mathbb{Z}/q^{I_{2}}\mathbb{Z})^{2}$.
Case 2: $q^{I_{3}}\geq(q^{I_{1}})^{\frac{1}{2}}$. In this case $\pi_{q,q^{I_{3}}}\circ\psi$ agrees with a homomorphism $h:G\rightarrow\Gamma_{q}\ltimes(\mathbb{Z}/(q^{I_{3}})^{\{\theta\}}\mathbb{Z})^{2}$ on a large subset $T$ of $G$, where $T$ boundedly generates a large subgroup $G^{\prime}$ of $G$. According to Proposition 3.24,
$\pi_{(q^{I_{3}})^{\{\frac{\theta}{4}\}},(q^{I_{3}})^{\{\frac{\theta}{4}\}}}(h(G))\subset\pi_{(q^{I_{3}})^{\{\frac{\theta}{4}\}},(q^{I_{3}})^{\{\frac{\theta}{4}\}}}(H_{u,v})$
for some $u,v\in\mathbb{Q}$.
Next, we apply Proposition 4.3 to find $(g_{0},w_{0})$ from $A\cdot A$ such that $g_{0}\in G^{\prime}$ and
$\pi_{p_{i}^{[\frac{1}{4}\theta n_{i}]}}(g_{0},w_{0})\not\in H_{u,v}(\operatorname{mod}p_{i}^{[\frac{1}{4}\theta n_{i}]}),\hskip 14.22636pt\forall i\in I_{3}^{\prime},$
where $I_{3}^{\prime}\subset I_{3}$ and $q_{3}^{\prime}\geq(q^{I_{3}})^{\frac{1}{2}}$.
On the other hand, since $T$ boundedly generates $G^{\prime}$, take $(g_{0},w_{1})\in(\psi(T))^{O(1)}\subset B^{O(1)}$. Let $(1,w^{\prime})=(g_{0},w_{0})(g_{0},w_{1})^{-1}$. Then
$w^{\prime}\not\equiv 0(\operatorname{mod}p_{j}^{[\frac{1}{4}\theta n_{j}]}),\forall j\in I_{3}^{\prime}.$
With $(1,w^{\prime})$ playing the same role as $(1,w)$ in the previous case, we can argue in the same way to construct a large subset of $\Gamma_{q}\ltimes(\mathbb{Z}/q_{3}^{\prime}\mathbb{Z})$.
In either case, we manage to construct a subset from a bounded product of $\{A\cup B\}$ whose projection to $\Gamma_{q}\ltimes(\mathbb{Z}/q^{\prime}\mathbb{Z})^{2}$ is very large for some large $q^{\prime}\|q$, starting with a set $B$ only known to have large projection to $\Gamma_{q}\ltimes\{\vec{0}\}$.
We can keep iterating the above argument to grow the modulus of the unipotent factor, and thus prove Proposition 2.2.